A body freely falling from a tower covers a distance of S=h/2 in the last second of it's motion.the height of tower is?

Grade:

3 Answers

akash luckey aggara
11 years ago

is the ans 10

shivang belwariar
11 years ago

we have the formula - S = U + a/2(2N-1) [ WHERE U = INITIAL SPEED , & N = value of TIME IN Tth SECOND . ]

so using the formula ------

 

h/2 = u + a/2(2t-1)

 

h/10 = 2t-1 { u=0 & a=10}

 

t = h+10/20

 

so, total time = h+10/20

now by formula = h=ut + 1/2at^2, (u=0 a=10)

 

h=h^2+100+20h/80

 

on solving - h = 58.28 metres.

 

plz like my answer

sharjil deshmukh
11 years ago

can we use \

S = u + g( total time - 1 -1/2}|

to get total time n thn subtitiute it in same eqation we wil get total distance|

is this ans right ???????????? 2 prsn who commented above